# LTLCrit: A Temporal Logic-Based LLM Critic for Safe and Efficient Embodied Agents

### v1

@@ Abstract
Large language models (LLMs) have demonstrated promise in reasoning tasks and general decision-making in static environments. In long-term planning tasks, however, errors tend to accumulate, often leading to unsafe or inefficient behavior, limiting their use in general-purpose settings. We propose a modular actor-critic architecture in which an LLM actor is guided by LTLCrit, a trajectory-level LLM critic that communicates via linear temporal logic (LTL). Our setup combines the reasoning strengths of language models with the guarantees of formal logic. The actor selects high-level actions from natural language observations, while the critic analyzes full trajectories and proposes new LTL constraints that shield the actor from future unsafe or inefficient behavior. The architecture supports both fixed, hand-specified safety constraints and adaptive, learned soft constraints that promote long-term efficiency. Our architecture is model-agnostic: any LLM-based planner can serve as the actor, and LTLCrit serves as a logic-generating wrapper. We formalize planning as graph traversal under symbolic constraints, allowing LTLCrit to analyze failed or suboptimal trajectories and generate new temporal logic rules that improve future behavior. We evaluate our system on the Minecraft diamond-mining benchmark, achieving 100% completion rates and improving efficiency compared to baseline LLM planners. Our results suggest that enabling LLMs to supervise each other through logic is a powerful and flexible paradigm for safe, generalizable decision making.
@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated strong performance on diverse reasoning and decision-making tasks, from natural language understanding, reasoning (Huang & Chang, 2022; Wei et al., 2022), and code generation (Li et al., 2022; Chen et al., 2021). However, much of this success has been in static, text-based settings. To deploy AI systems in real-world scenarios, it is essential to build agents capable of safe and efficient decision-making in complex, partially observable environments with long-term objectives and uncertainty.
Embodied environments such as Minecraft (Fan et al., 2022), Habitat (Savva et al., 2019), Alfworld (Shridhar et al., 2020) serve as useful testbeds for the next generation of agents. These environments require agents to perceive complex states, plan for long-horizon problems, under sparse supervision and rewards. Recent work has begun to explore the use of LLMs in embodied settings. SayCan Ahn et al. (2022) uses LLMs to decide affordance filtered actions for short-horizon robotic tasks, while Voyager (Wang et al., 2023a) leverages an LLM to autonomously explore and obtain as many new materials as possible in Minecraft.
@@
Despite these advances, LLMs continue to struggle with long-term sequential planning in embodied dynamical environments. While LLMs can often generate coherent short-term plans or respond to individual prompts, they lack the consistency, memory, and iterative refinement required for solving multi-step tasks where intermediate actions must align with long-term objectives. These challenges are especially pronounced in open-ended domains like robotics or Minecraft, where the agent must choose from a large set of possible actions while reasoning about tools, environment dynamics, and partial observability.
Recent evaluations (Kambhampati et al., 2024) show that LLMs struggle to reliably solve classical planning problems without the aid of external verification. Even simple domains expose brittleness: small prompt variations can cause cascading errors, and plans often violate task preconditions or logical constraints. These limitations have spurred the development of hybrid frameworks that pair LLMs with verifiers (Silver et al., 2022; Guo et al., 2024). While these methods improve correctness, they rely on hand-written task models.
@@
Ultimately, a central challenge is enabling foundational models like LLMs to act as reliable, autonomous agents in real-time, safety-critical, multi-agent environments such as healthcare (Hosny et al., 2018), automated transportation (Wang et al., 2023b), and domestic assistance (Birkmose et al., 2025). In each of these contexts, both safety and efficiency are crucial. Unsafe or inconsistent behavior can lead to failures in physical systems, and inefficient or incompetent agents rapidly lose the trust of human collaborators (Esterwood & Robert Jr, 2023). We argue that planning architectures should combine the reasoning strengths of LLMs with the formal structure of symbolic logic to ensure both flexibility and safety in real-world decision making.
@@ 1.1 Related works
Linear Temporal Logic (LTL) (Pnueli, 1977) is a formal language for specifying temporal properties of systems, for example, conditions that must hold at the next timestep given the current state, or conditions that must always hold, or eventually become true. LTL has been widely used in robot motion planning (Fainekos et al., 2005), for safe planning and control (Wongpiromsarn et al., 2012), and even in reinforcement learning Alshiekh et al. (2018). In each of these applications, LTL is used to specify safety constraints; however, one may also use it to specify efficiency constraints, as will be discussed in our graph formulation.
@@
Recent work has shown some promise in terms of LLMs’ abilities to perform in short-horizon planning tasks (Huang et al., 2022a). Particularly, Ahn et al. (2022) utilizes an LLM to asses the value of a given action, and weights it with an affordance function which quantifies the feasibility or success rate of the action given the current state of the system. It achieves a strong success rate in short-horizon mobile manipulation tasks. Huang et al. (2022b) introduces a closed-loop inner monologue by feeding language-based feedback back into the LLM during execution, enabling re-planning in dynamic environments. Simultaneously, other works such as Kambhampati et al. (2024) suggest that LLMs do not show reliability while operating autonomously, and require external critics. Our work builds directly upon these ideas, augmenting LLMs with formal verifiers. Finally, a majority of the work in LLM planning and policy learning talks about the success rate. Our emphasis is not only on completion rates, but also on speed of completion.
@@
There is also a growing body of work at the intersection of LLMs and temporal logic; however, the key emphasis is on using LLMs as tools for translation between natural language and LTL constraints Chen et al. (2023); Liu et al. (2023). Another line of work uses these translators in the loop (Yang et al., 2024), enforcing these constraints to filter out unsafe actions during execution. Compared to prior literature, which enables the use of static handcrafted rules, we allow an LLM critic to come up with further LTL constraints that show empirical efficiency gains in LLM planning problems.
@@
Recent work explores hybrid actor-critic setups where LLMs guide planning and evaluation, often using natural language feedback or prompt optimization loops (Dong et al., 2023; Yang et al., 2025). These models rely on soft, unstructured feedback and lack formal guarantees, particularly in multi-step embodied reasoning tasks. Our LTL-based trajectory critic offers interpretable, symbolic evaluation aligned with both safety and efficiency constraints.
@@ 1.2 Contributions
We propose LTLCrit, a novel Linear Temporal Logic-based Critic, which augments and boosts the performance of existing off-the-shelf LLM planners by protecting the planners against unsafe and inefficient actions. The composition of LTLCrit with an existing LLM planner leads to a novel symbolic actor-critic architecture designed to solve long-horizon planning tasks in dynamic, embodied environments using large language models (LLMs). Planning is decomposed across two timescales, a online actor loop that proposes high-level actions based on current state descriptions, and a offline critic loop that imposes symbolic constraints learned from past trajectories. Our modular decomposition aligns with LLMs’ strengths in local reasoning while addressing their weaknesses in long-term consistency over sequential domains.
LTLCrit is domain-agnostic, extensible to a variety of embodied tasks, and naturally integrates with existing LLM-based planners. It is the first to use symbolic temporal logic as a communication protocol between actor and critic modules, enabling interpretable, verifiable, and generalizable decision making. Our contributions are threefold.
@@
Symbolic Actor-Critic Architecture: We propose a two-timescale architecture where an LLM actor generates high-level actions in real time, and an LLM critic periodically analyzes trajectories to induce Linear Temporal Logic (LTL) constraints over symbolic abstractions of state and action. These constraints are used to prune unsafe or inefficient decisions, leading to improved sample efficiency and task performance. Our formulation allows constraints to be automatically discovered and continuously refined over time.
@@
Empirical Validation in Embodied Settings: We evaluate our method on challenging long-horizon tasks in Minecraft, where agents must autonomously mine a diamond. Our critic can improve both the consistency of behavior and average performance for two popular LLM planners, SayCan (Ahn et al., 2022) and InnerMonologue (Huang et al., 2022b).
Together, these contributions advance the goal of safe, efficient, and interpretable LLM-based decision-making in complex, dynamic environments. Our symbolic actor-critic framework bridges the gap between natural language reasoning and formal planning, opening up new possibilities for trustworthy embodied agents.
@@ 2 Background and Problem Formulation
Enabling LLM-based agents to work effectively alongside humans in embodied environments is a fundamental goal of LLM research. This has motivated recent efforts to test and enhance LLMs’ abilities in interactive, multi-step environments with long-term objectives. A central challenge LLM agents face in such settings is the lack of goal formalization and the difficulty of performing coherent long-horizon planning. Without structured guidance, LLM agents can behave reactively, inconsistently, or in an unsafe way. These behaviors are incompatible with reliable human-robot teaming.
Moreover, beyond safety, efficiency plays a critical role in collaboration. In human-agent teams, inefficient partners are perceived as unreliable, leading humans to reduce trust, ignore suggestions, or take full control (Fan et al., 2008). To deploy LLM agents in high-stakes or collaborative domains, we must ensure that their behavior is not only safe but also efficient, interpretable, and improvable over time.
In our work, we address these challenges by developing a learning mechanism that imposes formal guarantees on the behavior of LLM agents. Specifically, we aim to design agents that can self-correct and improve efficiency through structured feedback, from both humans and LLMs, using relatively little data. We study the planning problem in the context of Minecraft, an embodied environment, making LLM agents achieve the goal of mining a diamond.
@@ 2.1 Problem Statement
We formalize the problem of safe and efficient LLM planning in embodied environments as follows. Consider an agent operating in an environment with state space and a finite high-level action space , where . The agent observes a representation of the state, which is modeled by . Unlike standard RL settings, we assume no explicit reward function, leveraging instead the zero-shot learning capabilities of LLMs via a high-level goal description (e.g., “make an iron pickaxe” or “obtain a diamond”). We enforce safety and enable systematic constraint enforcement using an additional abstract boolean representation, , where , consisting of atomic propositions capturing important environmental features.
Within , there are two sets of interest.
@@
is the set of states where the high-level objective is satisfied.
@@
is the set of states that violate safety constraints.
We enforce behavioral constraints using Linear Temporal Logic (LTL) formulas over . We have two kinds of constraints.
Safety Constraints are expert-verified, hand-authored LTL constraints that prevent transitions into . These capture fundamental requirements such as collision avoidance.
Efficiency constraints are automatically generated by LTLCrit and eliminate suboptimal behaviors observed during execution, emerging from patterns such as redundant action sequences. For the sake of simplicity, we assume that every high level action is equally expensive to execute, and our goal is to minimize the number of primitives required to achieve the goal.
We seek to develop an agent that satisfies all safety constraints with high probability and efficiently performs a sequence of actions that leads to goal states.
2.2 Linear Temporal Logic Primer

### v2

# LogicGuard: Improving embodied LLM agents through temporal logic based critics

@@ Abstract
Large language models (LLMs) have shown promise in zero-shot and single step reasoning and decision-making problems, but in long-horizon sequential planning tasks, their errors compound, often leading to unreliable or inefficient behavior. We introduce LogicGuard, a modular actor–critic architecture in which an LLM actor is guided by a trajectory-level LLM critic that communicates through Linear Temporal Logic (LTL). Our setup combines the reasoning strengths of language models with the guarantees of formal logic. The actor selects high-level actions from natural language observations, while the critic analyzes full trajectories and proposes new LTL constraints that shield the actor from future unsafe or inefficient behavior. LogicGuard supports both fixed safety rules and adaptive, learned constraints, and is model-agnostic: any LLM-based planner can serve as the actor, with LogicGuard acting as a logic-generating wrapper. We formalize planning as graph traversal under symbolic constraints, allowing LogicGuard to analyze failed or suboptimal trajectories and generate new temporal logic rules that improve future behavior. To demonstrate generality, we evaluate LogicGuard across two distinct settings: short-horizon general tasks and long-horizon specialist tasks. On the Behavior benchmark of 100 household tasks, LogicGuard increases task completion rates by 25% over a baseline InnerMonologue planner. On the Minecraft diamond-mining task, which is long-horizon and requires multiple interdependent subgoals, LogicGuard improves both efficiency and safety compared to SayCan and InnerMonologue. These results show that enabling LLMs to supervise each other through temporal logic yields more reliable, efficient and safe decision-making for both embodied agents.
@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated strong performance on diverse reasoning and decision-making tasks, from natural language understanding, reasoning (Huang & Chang, 2022; Wei et al., 2022), and code generation (Li et al., 2022b; Chen et al., 2021). However, much of this success has been in static, text-based settings. In contrast, embodied dynamical environments require agents to plan over long horizons under uncertainty, partial observability, and complex dynamics. While LLMs can generate short-term plans or respond coherently to individual prompts, they lack the consistency, memory, and iterative refinement needed to solve multi-step tasks where intermediate actions must align with long-term goals. These shortcomings are especially pronounced in open-ended domains such as robotics and interactive virtual environments, where agents must reason over large action spaces, tools, and environment dynamics.
Recent evaluations (Kambhampati et al., 2024) show that even in simplified settings, LLMs often fail to produce reliable plans without external verification. Small prompt variations lead to compounding errors, and generated plans frequently violate preconditions or logical constraints. To address these weaknesses, hybrid frameworks pair LLMs with external verifiers (Silver et al., 2022; Guo et al., 2024). Yet, these methods typically rely on manual design and labeling, limiting scalability. In this work, we aim to reduce such manual effort by automatically generating formal constraints that guide and safeguard LLM planning.
Embodied environments such as Minecraft (Fan et al., 2022; Wang et al., 2023a), iGibson (Li et al., 2022a), and VirtualHome (Puig et al., 2018) serve as useful testbeds for developing such architectures, while datasets like Behavior (Li et al., 2023) benchmark agent performance across diverse tasks. These domains capture core challenges of embodied intelligence, perceiving high-dimensional states, planning under sparse supervision, and executing long-horizon strategies. Recent work has begun to apply LLMs in these settings, from household robotics (Ahn et al., 2022) to resource-driven virtual worlds (Wang et al., 2023a), but scalability and reliability remain open problems.
@@
Ultimately, enabling LLMs to function as trustworthy autonomous agents requires robustness in safety-critical, long-horizon, and multi-agent contexts such as healthcare (Hosny et al., 2018), automated transportation (Wang et al., 2023b), and domestic assistance (Birkmose et al., 2025). In these domains, unsafe or inconsistent behavior risks physical failures, while inefficiency erodes human trust (Esterwood & Robert Jr, 2023). We argue that planning architectures must combine the flexible reasoning of LLMs with the formal guarantees of symbolic logic. To this end, we propose a symbolic actor-critic framework that uses temporal logic to enforce safety, improve performance, and ensure interpretable decision-making in embodied environments.
@@ 1.1 Related works
Linear Temporal Logic (LTL) (Pnueli, 1977) is a formal language for specifying temporal properties of systems through boolean logic based expressions. LTL has been widely used in robot motion planning (Fainekos et al., 2005), for safe planning and control (Wongpiromsarn et al., 2012), and even in reinforcement learning (Alshiekh et al., 2018). In each of these applications, LTL is used to specify safety constraints; instead, we shall use it to specify performance constraints.
@@
Recent work has explored LLMs for short-horizon planning (Huang et al., 2022a). SayCan (Ahn et al., 2022) scores actions with an LLM and weights them by an affordance function, while InnerMonologue (Huang et al., 2022b) feeds LLM feedback back into itself to enable online re-planning. In embodied environments, Ziliotto et al. (2025) uses compositional planning for LLMs, while Wu et al. (2023) uses alignment to improve reasoning in embodied environments. Chen et al. (2024) breaks down complicated tasks into subgoals. Other works such as (Kambhampati et al., 2024) suggest that LLMs do not show reliability while operating autonomously, and require external critics. Our work builds directly upon these ideas, augmenting LLMs with formal verifiers.
@@
Recent work at the intersection of LLMs and temporal logic focuses on translating natural language into LTL constraints (Chen et al., 2023; Liu et al., 2023), or enforcing such translations during execution to filter unsafe actions (Yang et al., 2024). These methods use static handcrafted rules. Recent work (Ravichandran et al., 2025) uses LLMs to generate safety constraints online for robotics, their focus is purely on avoiding unsafe behaviors. In contrast, we leverage an LLM-based LTL law generator not only to ensure safety but also to actively improve planning performance and long-horizon task performance, enabling adaptive, empirically grounded constraint generation in complex sequential environments.
@@
Prior literature explores hybrid actor-critic setups where LLMs guide planning and evaluation, using natural language feedback or prompt optimization loops (Dong et al., 2023; Yang et al., 2025). These models rely on unstructured feedback and lack formal guarantees, particularly in multi-step embodied reasoning tasks. Our LTL-based trajectory critic offers interpretable, symbolic evaluation enforcing both safety and performance constraints.
@@ 1.2 Contributions
We propose LogicGuard, a novel Temporal Logic-based critic, which augments and boosts the performance of existing off-the-shelf LLM planners by protecting them against unsafe and inefficient actions. Composing LogicGuard with an LLM planner leads to a novel symbolic actor-critic architecture designed to solve long-horizon planning tasks in dynamic, embodied environments using large language models (LLMs). This architecture breaks sequential planning into two timescales; an online actor proposes high-level actions based on current state descriptions, and an offline critic loop that imposes symbolic performance and safety constraints learned from past trajectories. This modular decomposition leverages LLMs’ strengths in local reasoning and addresses their weaknesses in long-term consistency.
LogicGuard is domain-agnostic, naturally integrates with existing LLM-based planners, and is the first to use symbolic temporal logic as a communication protocol between the actor and critic, enabling interpretable, verifiable, and generalizable decision making. Our contributions are threefold.
@@
Symbolic actor-critic architecture: We propose a two-timescale architecture where an LLM actor generates high-level actions online, and an LLM critic periodically analyzes trajectories offline to induce Linear Temporal Logic (LTL) constraints on the actor. These constraints are meant to prune unsafe or inefficient decisions, improving task performance. Our formulation allows constraints to be automatically discovered and continuously refined over time.
@@
Empirical validation in generalist and specialist embodied settings: We demonstrate gains in (i) completion rates on 100 short-horizon household tasks from the Behavior dataset, where atomic propositions are automatically derived across diverse environments, and (ii) efficiency and consistency in the long-horizon Minecraft diamond-mining task, where hand-engineered propositions capture domain structure. This demonstrates that our architecture benefits both generalist agents operating in varied environments and specialist agents in structured long-horizon domains.
Together, these contributions advance the goal of robust, safe, and interpretable LLM-based decision making in complex, dynamic environments through self-supervision, bridging the gap between natural language reasoning and formal planning.
@@ 2 Background and Problem Formulation
A key challenge in deploying LLM-based agents in embodied environments is their difficulty with coherent long-horizon planning. Without explicit goal formalization or structured guidance, they often act reactively or inconsistently, leading to unsafe and inefficient behavior that undermines reliability in human-agent teams (Fan et al., 2008). For real-world collaboration, agents must not only avoid unsafe actions but also demonstrate efficiency, competence, and interpretable reasoning.
We address this by developing a mechanism that provides formal guarantees on agent behavior while allowing for self-correction and improvement from structured feedback. Our framework uses symbolic constraints learned over time to guide LLM decision-making, ensuring safe and efficient behavior with limited data. We evaluate this approach across two settings: (i) short-horizon household tasks in the Behavior dataset, and (ii) the long-horizon challenge of mining a diamond in Minecraft.
@@ 2.1 Problem Statement
We formalize the problem of safe and efficient LLM planning in embodied environments. Consider an agent operating in a state space and a finite high-level action space , where . The agent observes a representation of the state, which is modeled by . Unlike standard RL, we assume no explicit reward function, and instead rely on a natural language goal description (e.g., “make an iron pickaxe” or “obtain a diamond”).
To enable formal reasoning, we introduce an abstract Boolean representation , where , encodes atomic propositions capturing salient environment features. Within , we define , the set of states that satisfy task objectives, and , the set of states that violate safety requirements.
Behavior is regulated through two types of LTL constraints. Safety constraints are expert-authored rules that forbid transitions into , such as collision avoidance. Performance constraints, in contrast, are automatically induced by LogicGuard from observed trajectories, eliminating redundant or suboptimal action patterns. Assuming all high-level actions have equal cost, the agent’s objective is to minimize the number of primitive actions required to reach while satisfying all safety constraints with high probability.
2.2 Linear Temporal Logic Primer
Linear Temporal Logic(LTL) (Pnueli, 1977) provides a formal language for expressing temporal properties over sequences of states. LTL formulas consist of variables called atomic propositions, Boolean operators, and temporal operators(discussed in Appendix A.1). LTL formulas can be converted into Büchi automata, which are finite-state machines enabling the algorithmic verification of whether a trajectory satisfies a given temporal specification. Using atomic propositions corresponding to physically meaningful events leads to interpretability. We employ SPOT (Duret-Lutz & Poitrenaud, 2004) to enforce LTL laws on the LLM actor.
2.3 Experimental Domains
We evaluate LogicGuard in two contrasting embodied environments. Behavior (Li et al., 2023) consists of short-horizon household tasks in diverse environments, with multiple independent subtasks. In contrast, Minecraft presents long-horizon compositional tasks with interdependent subgoals; we study the diamond-mining task (Guss et al., 2019). These domains allow us to evaluate LogicGuard on both generalist agents in diverse, short-horizon tasks and specialist agents in structured, long-horizon tasks.
3 Methods
3.1 Architecture Overview
We propose a hierarchical planning architecture that integrates large language models (LLMs) with formal symbolic reasoning to achieve safe and efficient decision-making in complex environments. The architecture consists of two interacting loops operating at different timescales:
Online actor loop:
At each timestep, an LLM actor is given a natural language state description, and chooses a high-level action . The action is verified against current LTL constraints, defined over . If valid, it is executed via a low level controller. Otherwise, the actor is informed that the chosen action is invalid and prompted to choose a new action. Examples of the LTL verifier in action are shown in Figure 2.
Offline critic loop:
Periodically, an LLM-based critic analyzes completed trajectories to identify incorrect, inefficient or unsafe behaviors, proposing new LTL constraints or removing existing ones. Updates are immediately incorporated into the verifier, affecting subsequent actions.
The separation of online reactive planning and offline symbolic refinement enables safe, efficient and interpretable decision-making. The modular design also supports easy transfer across domains. A full architectural diagram is presented in Figure 1.
3.2 Defining the atomic propositions
The choice of variables in directly affects the critic’s expressive power and tractability. For generalist environments like Behavior, we automate the design of , initializing it with variables needed for goal satisfaction and safety constraints, and then augmenting it based on observed state changes during exploratory rollouts. This ensures the abstract state is both task-aware and environment-adaptive while avoiding excessive manual design (details in Appendix A.2.1).
3.3 Planning as a Graph Traversal Problem
We frame efficient planning as a shortest path problem under safety constraints. Each primitive action has unit cost, as LLM calls dominate execution time. The agent aims to reach a state in from its current state in as few steps as possible, while avoiding .
We model the problem as a bipartite graph . One partition of this graph consists of symbolic states , while the other partition consists of . Edges from to exist only if is allowed by current LTL constraints; edges from to represent state transitions. This bipartite structure explicitly separates the roles of the actor, which selects edges from states to state-action pairs, and the critic, which prunes edges via constraints.
Due to an exponential number of nodes, finding the shortest path in this graph is exponentially hard, motivating the need for LLMs to guide exploration via natural language reasoning, efficiently navigating the graph despite its combinatorial complexity.
3.4 LLM Actor: Guided Exploration
The LLM actor receives a full state description and proposes a high-level action from the action space such as “mine_stone” or “grasp_plywood”. In the bipartite graph , the actor chooses an edge flowing out from the agents current state , which is legal as per the current LTL constraints. The use of the LLM allows us to replace brute force exploration with semantically guided traversal in a large symbolic space. Our architecture allows the direct use of existing LLM planners. We adopt InnerMonologue (Huang et al., 2022b) for Behavior, and both SayCan (Ahn et al., 2022) and InnerMonologue for Minecraft. To ensure adaptivity and prevent overly restrictive rules, the actor tracks repeated attempts of violations of LTL constraints; if a particular rule is triggered beyond a fixed threshold, it is removed, allowing the agent to explore alternative strategies.
3.5 LogicGuard: the Linear Temporal Logic-based LLM Critic
In our experiments, we choose to run the critic to analyze complete trajectories. In practice the critic may be run more or less often depending on the requirements of the specific application. The critic is prompted to identify inefficient behaviors and to propose new LTL constraints of the form
where is a boolean expression over symbolic state features, and specifies allowed or disallowed actions. These constraints eliminate inefficient behaviors. For example,
prevents crafting duplicate wooden pickaxes. Constraints are only generated if is observed in the trajectory, ensuring generalization grounded in data. Constraints are induced from three sources:
@@ 3.5 LogicGuard: the Linear Temporal Logic-based LLM Critic
Environment feedback: Actions deemed invalid by the environment (e.g. soaking a rag without turning on the tap or mining diamonds without an iron pickaxe) lead to an error message from the environment. These actions are encoded into constraints to prevent future errors.
@@
Graph-based efficiency: The critic is prompted to analyze the symbolic task graph from 3.3. The critic is asked to classify actions as efficient or inefficient, pruning wasteful actions, with an emphasis on repetitive actions.
3.
Overconstrained States: When current laws collectively eliminate all feasible actions in a state, the system falls back to bare minimum hand-engineered laws. Offline, these states are analyzed to refine or relax constraints, preventing overly restrictive rules.
Since the critic is only allowed to induce constraints grounded in trajectory data, all proposed rules are traceable, and we may bound the number of possible rules that are generated.
Theorem 1.
Let be a trajectory with , and each a one-hot vector representing an action from a finite action space of size . Consider an algorithm that generates LTL constraints of the form equation 1 where is a boolean condition over the symbolic state that holds for at least one in the trajectory and .
Proof.
There are at most unique values for . Each constraint is of the form , and must hold for at least one , the number of distinct that can be constructed from the trajectory is at most . Further, the algorithm can only decide if each action is allowed or disallowed. As a consequence, the total number of laws is at most . However, since actions cannot be simultaneously allowed and disallowed, the number of actions an algorithm operating under our assumptions can make such that every state has at least one feasible action is at most . ∎
In our bipartite graph representation, pruning edges based on observed trajectories reduces complexity from to linear in dataset size. Interpretable atomic propositions allow the critic to generalize constraints to semantically equivalent but unseen states. Finally, the sparse structure of goal-directed task graphs allows LogicGuard produces sample-efficient and robust behavior while enforcing both safety and efficiency.
4 Experiments
4.1 Experimental Setup
Our actors use OpenAI GPT-4.1 as a backbone LLM and our critics use o3-mini. All temperatures are set to 0.1 to reduce stochasticity while still allowing exploration. As discussed in Section 2.3, we evaluate LogicGuard in two embodied environments. Implementation details including prompts and APs are presented in the appendix.
Behavior:
We use the Behavior (Li et al., 2023) dataset, consisting of short-horizon (average horizon: 14.6) household tasks with multiple independent subtasks. High-level actions and observations are designed via the API and goal specifications from Li et al. (2024). Completion rate is the primary metric, as prior work shows non–chain-of-thought LLMs struggle in this domain. To support diverse tasks and environments, atomic propositions are automatically generated, enabling LogicGuard for general diverse settings. We adopt InnerMonologue as the LLM actor. Given the diversity of tasks and environments, implementing a reliable affordance function is challenging, which limits the applicability of SayCan in this setting.
Minecraft:
Minecraft provides a partially observable environment with compositional, interdependent subgoals. We study the diamond-mining task (Guss et al., 2019), a long-horizon setting (see Table 2) requiring a sequence of dependent subgoals without intermediate rewards. We interface via the Mineflayer API (PrismarineJS Team, 2025), which provides structured observations and path-planning utilities, upon which we design atomic actions. Evaluation metrics in this domain are efficiency (number of high-level actions required to obtain a diamond) and safety (number of failed or illegal actions). Unlike in Behavior, here we design hand-engineered atomic propositions, highlighting LogicGuard’s ability to augment specialist agents for complex compositional goals.
4.2 Baselines
Our modular architecture allows us to augment off-the-shelf LLM planners with our symbolic critic. We focus on two representative planners:
InnerMonologue (Huang et al., 2022b):
An LLM planner that interleaves natural language thoughts and code actions, incorporating feedback at each step. This provides implicit reflection and sequential planning. We use InnerMonologue in both Behavior and Minecraft.
SayCan (Ahn et al., 2022)
A two-stage planner that filters feasible actions via affordances and ranks them for goal relevance. We only evaluate SayCan in Minecraft, since affordance functions do not scale to Behavior’s large, diverse action space. In Minecraft, SayCan is combined with LogicGuard for LTL-based affordance filtering.
In Minecraft, LogicGuard refines constraints for both InnerMonologue and SayCan until performance stabilizes (two iterations sufficed). In Behavior, LogicGuard augments InnerMonologue only, and the critic is invoked on failed tasks iteratively for up to two iterations to prevent overconstraining successful trajectories. While LogicGuard supports hand-engineered LTL constraints, all our experiments use only critic-generated laws. The only exception is the SayCan baseline in Minecraft, which requires a hand-engineered affordance function.
4.3 Results
4.3.1 Behavior Benchmark: Task Completion
Behavior (Li et al., 2023) consists of 100 short-horizon household tasks with multiple independent subtasks (e.g., pick up objects, place items, open containers). Since tasks are short-horizon, we focus on task completion rather than efficiency. We end all trajectories after 40 actions, or if the actor chooses to declare it is done.
| Method | Completed Tasks |
| InnerMonologue | 47% |
| InnerMonologue + LogicGuard | 72% |
4.3.2 Minecraft: Efficiency and Safety
In Minecraft, the agent must mine a diamond from scratch, involving long-horizon dependencies across mining and crafting subgoals. We evaluate efficiency (number of primitive actions to reach key subgoals) and safety (number of failed or unsafe actions).
Efficiency
Table 2 shows the average number of primitives required to reach each subgoal. We note a 23% increment in the performance of InnerMonologue in the diamond mining task. SayCan is very easily distracted by the abstract nature of high level actions such as “explore”, and does not make meaningful progress on the task. Our architecture identifies these drawbacks and blocks exploration related actions till the right tools are available.
| Method | Wooden Tool | Stone tool | Iron Tool | Diamond |
| SayCan | N/A (0/5) | N/A (0/5) | N/A (0/5) | N/A (0/5) |
| SayCan + LogicGuard | 12.6(5/5) | 17.6(5/5) | 37.8(5/5) | 45.4(5/5) |
| InnerMonologue | 12.2 (5/5) | 18.2 (5/5) | 43.25 (4/5) | 45.5 (4/5) |
| InnerMonologue + LogicGuard | 9.4 (5/5) | 14.4 (5/5) | 32.0 (5/5) | 35.8 (5/5) |
Safety
We measure failed actions and the number of unsafe actions blocked by the critic. LogicGuard significantly reduces both failure rates and unsafe actions (Table 3). Failures after LogicGuard are due to path planning errors within Mineshafter’s API.
| Method | Failed Actions | Critic-Blocked Unsafe Actions |
| InnerMonologue | 23% | N/A |
| InnerMonologue + LogicGuard | 4.5% | 15% |
Our results highlight that LogicGuard generalizes across task structure and horizon, helping LLM actors act more reliably, safely, and efficiently.
5 Discussion
5.1 LogicGuard mitigates systematic LLM actor failures
Naive LLM actors often repeat actions after task completion or ignore environment feedback, leading to inefficiency and loops. In Behavior, an actor may repeatedly pick up or place already organized objects; in Minecraft, it can mine blocks beyond task requirements. These failures stem from limited reasoning and prompt overload. LogicGuard addresses these issues by detecting redundant or failed actions and blocking them based on task conditions. This enforces interpretable, verifiable constraints, breaking loops and improving both reliability and task efficiency.
5.2 Atomic propositions are a key design choice
Atomic propositions (APs) define the variables the critic uses to construct LTL constraints, directly controlling rule expressivity. In the Behavior dataset, the most common failure mode for LogicGuard is because our APs are insufficiently expressive. For instance, our current automated AP generation treats each item in the environment as an independent entity, which complicates combinatorial tasks. Consider placing four plates into four boxes such that each box contains at least one plate. There are 24 feasible solutions. Encoding a single LTL formula that captures all feasible final states is highly complex for the critic, which must account for all possible permutations. By contrast, in Minecraft we hand engineered APs to precisely capture task-relevant state features. This targeted design simplifies rule generation, reduces combinatorial complexity, and enables more reliable constraint synthesis.
5.3 Generalist vs Specialist agents
LogicGuard improves performance across diverse domains. In Behavior, the environment is unknown and contains rules the LLM cannot anticipate (e.g., items must be inside a sink before soaking), which the critic must discover iteratively, akin to human learning. In Minecraft, extensive textual documentation allows the LLM to succeed with minimal guidance. Evaluating both domains demonstrates that LogicGuard supports generalist agents in novel settings and specialist agents in structured, long-horizon tasks.
6 Conclusion and Future Work
We introduced a modular actor-critic architecture in which an LLM critic supervises an LLM actor using linear temporal logic (LTL) constraints over abstracted versions of full trajectories. The critic operates at a slower timescale and applies zero-shot reasoning to identify and correct unsafe or inefficient behavior in a few iterations. Our modular architecture enables us to use existing off-the-shelf LLM planners as actors. The use of LTL constraints guarantees shielding of the LLM from unsafe or inefficient behavior. By expressing constraints in LTL over human-readable atomic propositions, we gain a symbolic structure that is immediately enforceable and human verifiable. We demonstrate our approach in two contrasting embodied domains, achieving significant improvements over baseline off-the-shelf LLM agents.
Several directions remain for future work. First, a critic could leverage annotated expert trajectories to imitate optimal behavior. Second, in dynamic environments, incorporating mathematical models of the environment into the critic could enable generation of LTL laws that evolve over time. Such adaptive laws are particularly relevant for multi-agent coordination and human-robot teaming, where modeling other agents can inform constraint synthesis. Finally, we aim to extend these ideas to real-world robotics, where online decisions may rely on smaller, potentially unreliable models that require formal constraints. Overall, our approach demonstrates that formal logic-based constraints provide a promising path toward safe, scalable, and general-purpose LLM agents.
7 Ethics Statement
Our work focuses on improving LLM-based planners in simulated environments (Behavior and Minecraft) and does not involve human subjects or sensitive data. LLMs can exhibit unsafe or unreliable behavior, and overreliance on them in real-world robotics could be hazardous. Our research hopes to draw attention and inspire further work towards safety nets for blackbox foundational models. Our framework is intended as a first step to improve reliability and safety in LLM-driven agents.
8 Reproducibility statement
We provide detailed descriptions of all environments, prompts and LLM models used in our experiments. While we use OpenAI’s GPT-4.1 and o3-mini APIs, which are inherently stochastic even at low temperature.
Acknowledgments
We acknowledge the use of ChatGPT for assistance in improving the wording and grammar of this document.
Appendix A Appendix
A.1 Linear Temporal Logic Operators